OpenStudy (anonymous):

(02.06) Which graph represents the solutions to the inequality |2x - 6| less than or greater to 10? number line with a closed circle on negative 2, shading to the right and a closed circle on 8, shading to the left number line with a closed circle on negative 2, shading to the left and a closed circle on 8, shading to the right number line with an open circle on negative 2, shading to the right and an open circle on 8, shading to the left number line with an open circle on negative 2, shading to the left and an open circle on 8, shading to the right

OpenStudy (anonymous):

i'll give medal

OpenStudy (anonymous):

\[\left| 2x-6 \right|\le10\] -> \[2x-6\le10\]\[x \le 8\] \[-(2x-6) \le 10\]\[-2x \le 4\]\[x \ge -2\] (remember to switch the direction when dividing or multiplying by a negative) Therefore you have closed circles and it is between -2 and 8!

OpenStudy (anonymous):

The circles are closed because it is less than or EQUAL TO. Key words that you can use to throw out the last two answers right away.